Calculate the force of gravity between a newborn baby of mass 4 kg and the obstetrician of mass 75 kg, who is 0.2 m from the baby.
Which exerts more gravitational force on the baby, Mars or the obstetrician, given that Mars has a mass of 6.4 1023 kg and is 8. 1010 m from Earth?
1
Expert's answer
2012-10-30T10:59:13-0400
Force between a baby and the obstetrician is
F = G*m1*m2/r^2 = 6.67300*10^{-11} *4*75/(0.2^2) =5*10^{-7} N
Force between a baby and Mars is
F = 6.67300*10^{-11}*4*6.4*10^{23}/((8*10^10)^2)= 2.67*10^{-9} N
obstetrician exerts more gravitational forcethen Mars.